When a child breaks open her piggy bank, she finds a total of 46 coins, consisting of nickels, dimes, and quarters. The total value of the coins is dimes were nickels, the value of the coins would be $4. How many nickels, dimes, and quarters were in the piggy bank? nickels dimes quarters Talk to a Tutor Need Help? İleti liene inene Watch It My Notes Ask Your Teacher O
Answer #1

Let

nickels =x

dimes =y

quarters =z

Since total coins is 46

x+y+z = 46

since total value is 5,

0.05x +0.1y +0.25z = 5

next

0.10x +0.05y+0.25z =4

So, we have system of equation

x+y+z = 46

0.05x +0.1y +0.25z = 5

0.10x +0.05y+0.25z =4

Solving we get

x=10

y= 30

z =6

hence

nickels = 10

dimes = 30

quarters =6